# SDPO: Segment-Level Direct Preference Optimization for Social Agents

### v1

@@ 4.3 Baselines
We compare the proposed SDPO with several strong baselines. 1) OpenAI proprietary LLMs. We provide the specific model versions in Appendix D.1. 2) SFT Behavioral Cloning fine-tunes LLMs on expert interaction data, producing a resulting model that serves as the base agent for SDPO and the following baselines. 3) DPO optimizes the agent policy based on turn-level data, specifically targeting the first differing turn in the positive and negative samples used by SDPO. 4) ETO optimizes the agent policy using session-level data. ETO utilizes the same negative sessions as SDPO while sampling five new sessions from scratch to form the data pairs. 5) DMPO leverages the same data as ETO and employs a new loss function to update the policy. 6) Preferred-SFT fine-tunes the base gent on the positive sessions in SDPO.
@@ 4.4 Results
Figure 3: The goal ratings and average words per session for various agents. The word count includes only the utterances of our agents and excludes those of GPT-4o. The square bracket denotes [average turns per session, average words per turn].
We present the results of SDPO and all the baselines on SOTOPIA in Table 1. As shown, in both the goal and relationship dimensions, SDPO significantly outperforms turn-level DPO, session-level ETO, and DMPO, even surpassing proprietary LLMs like GPT-4o by a large margin, highlighting the effectiveness of segment-level alignment. By analyzing the interaction histories in SOTOPIA, we find that weaker agents often exhibit stubbornness and only express their demands repeatedly. This leads to lower goal and relationship levels, especially in self-chat scenarios. Behavioral cloning using expert data can effectively improve this situation, making the agent more communicative. The reason why Llama-8B+BC’s goal rate drops in its interaction with GPT-4o is that the agent becomes persuadable. We also observe that aligned agents achieve simultaneous improvements in both goal and relationship. This indicates that alignment methods indeed enhance the social intelligence of models, rather than achieving goals through behaviors that violate social norms like threatening and deception.
We also repeat the above experiments using Mistral-Instruct-v0.3 Jiang et al. (2023), with the results presented in Table 2. The detailed experimental setup for Mistral is provided in Appendix D.2. SDPO consistently outperforms all baselines, demonstrating the generalization of our method.
@@ 4.5 Analysis
Necessity of Multi-turn Alignment After DPO adjusts the first-turn output probabilities for positive and negative segments, will the probabilities of positive segments increase and those of negative segments decrease in subsequent turns? To explore this, we plot the probability differences between positive and negative segments for DPO and SDPO during training, as shown in Figure 2 (only SDPO can be directly compared with DPO; therefore, ETO and DMPO are not mentioned here.). The DPO-turn trajectory is nearly parallel to the DPO trajectory, indicating that DPO has almost no influence on the probability differences of subsequent turns. In contrast, the SDPO trajectory rises more steeply. These results demonstrate the necessity of explicitly modifying the probability distribution across turns within the entire segment, providing an explanation for the superiority of multi-turn alignment over DPO.
Variation in Model Output Length We present the output length of various agents during their interactions with GPT-4o in Figure 3. Compared to the BC agent, all alignment methods increase the output length of the models. This phenomenon is commonly observed when DPO is applied to AI chatbots. However, unlike the users’ potential bias toward longer responses, which might be misleading, effective social strategies in social scenarios often require more tokens for communication. Thus, the increase in output length is reasonable. Furthermore, we experiment with terminating the dialogue when the SDPO-tuned agent reaches 8 interaction turns. This approach achieves better performance using fewer tokens than LLama-3.1-8B, demonstrating that the SDPO-tuned agent utilizes tokens more efficiently.
Efficiency of Positive Sample Utilization The quality of positive sessions sampled at the session level and segment level is illustrated in Figure 4. When the sampling count is consistent, session-level positives outperform segment-level ones in both goal and relationship. Sampling from scratch provides a larger sampling space compared to using partial interaction history, increasing the likelihood of generating high-quality sessions. However, some high-scoring sessions result from behavioral changes by the other agent, which prevents the model from learning correct behavioral patterns. In SDPO, the interaction history prior to erroneous turns introduces constraints to the action space of the interlocutor, effectively reducing such occurrences. Consequently, SDPO makes better use of positive samples and achieves superior performance compared to ETO and DMPO.
@@
4.6 Ablation Study
@@
$[m,n]$ 8.19 3.34 7.97 3.01
Table 3: Performance comparison of various segment selection methods. $A_{w}$ denotes the length of the segment selected by GPT-4o from each positive sample, and its distribution is presented in Appendix C.1. $m,n$ represent the sums of all turns in negative and positive samples, respectively, after the erroneous rounds.
Segment Selection We explore different segment selection methods of SDPO, with the results presented in Table 3 (in the square brackets, the length of the negative segment is listed first, followed by the positive segment.). The segment length refers to the number of turns contained within the segment. For symmetric segment lengths, methods with fixed lengths of 3 and 5 outperform the length of 1 (DPO), demonstrating the efficacy of multi-turn alignment. The method with a segment length of 5 is less effective than that with a length of 3, indicating that longer segments are not always better. Building on this insight, we leverage GPT-4o to automatically identify key segments from each positive sample, achieving the best results. For asymmetric segment lengths, model training for segment lengths of [3,1] and [5,3] collapse and can not interact normally. Other asymmetric segments underperform compared to their symmetric counterparts, supporting the theoretical discussions in Sections 2.3 and 3.3. Furthermore, we observe that as the degree of asymmetry decreases, the model’s performance improves. This improvement could be attributed to the reduced effect caused by the uneliminated $Z$ on the loss as asymmetry diminishes. This finding also helps explain the effectiveness of ETO, which does not impose constraints on the lengths of positive and negative sessions.
@@
Table 4: Performance comparison of models trained on data from different sources. The "Source" column indicates interaction participants: [test agent | interlocutor].
@@ 5 Related Work
Alignment Methods of Different Granularities Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022) is an effective approach for aligning LLMs’ outputs with human preferences. However, it has notable drawbacks, including instability, complexity, and high resource consumption. To address these challenges, Rafailov et al. (2023) introduce DPO, which simplifies training by using offline data with a straightforward classification loss. We classify DPO as a turn-level alignment method, and various alignment algorithms at different granularities have been developed based on DPO. Token-level DPO (TDPO) Zeng et al. (2024) integrates forward KL divergence constraints at the token level, enhancing both alignment and diversity. Step-DPO Lai et al. (2024) utilizes individual reasoning steps for preference optimization instead of holistic answer-level evaluation. However, in multi-turn interaction scenarios such as goal-oriented social dialogues or web navigation, single-turn alignment is insufficient. To tackle this, ETO and DMPO extend DPO to multi-turn contexts by leveraging session-level data. We take a step further by proposing SDPO, which introduces a segment-level sample pair optimization framework to achieve finer-grained alignment in multi-turn interactions.
@@ 6 Conclusion
In this paper, we introduce Segment-Level Direct Preference Optimization (SDPO) to improve the performance of LLM-based agents in multi-turn, goal-oriented social dialogues. Unlike existing alignment methods such as turn-level DPO and session-level approaches including ETO and DMPO, SDPO focuses on optimizing the agent policy by targeting specific key segments within a session. Our extensive evaluation on the SOTOPIA benchmark shows that SDPO significantly outperforms existing methods, highlighting the superiority of segment-level alignment. Looking ahead, we plan to apply SDPO to other agent tasks to further explore its versatility and effectiveness.
@@ Appendix B Supplementary Theoretical Analysis
B.1 DPO
Rafailov et al. (2023) propose Direct Preference Optimization (DPO), a method that leverages pairwise preference data to train policy models. In the context of social dialogue, we denote the number of the erroneous round as $e$. The DPO loss function is as follows:
where $y_{e}^{w},y_{e}^{l}\sim\pi_{\theta}(\cdot|h_{e})$ represent positive and negative output in the erroneous turn respectively.
B.2 ETO
Song et al. (2024) propose Exploration-Based Trajectory Optimization (ETO), which extends turn-level DPO to the session level without rigorous proof. The loss function is as follows:
@@ B.2 ETO
where $h^{w}$, $h^{l}$ represent complete positive and negative interaction histories respectively, $T_{w},T_{l}$ denote the number of rounds in each.
B.3 Discussion on DMPO
In Eq (5), the reward for the entire sequence should be calculated as the summation over all $(s,a)$ pairs. Let’s first discuss why it is valid to sum over time steps $t$. For LLMs, the state $s$ can be viewed as the input context, while $a$ represents the model’s output. In multi-turn interactions, all $(s,a)$ pairs within the sequence are unique. Thus, summing over time steps t is equivalent to summing over $(s,a)$ pairs, making the process more straightforward. However, essentially, each $(s,a)$ pair should be treated equally in the summation, without any inherent concept of time step $t$. Therefore, introducing a discount factor $\gamma^{t}$ is not appropriate. The correct Eq (5) is shown as follows:
$p(\tau^{w}\succ\tau^{l}|s_{0})=\\ \sigma\left(\sum_{t=0}^{T_{w}-1}r(s_{t}^{w},a_{t}^{w})-\sum_{t=0}^{T_{l}-1}r(s_{t}^{l},a_{t}^{l})\right).$ (10)
Moreover, DMPO introduces regularization for rounds based on Eq (5) to eliminate $Z$. However, it does not discuss the impact of length normalization, and this transformation lacks rigorous theoretical justification.

### v2

@@ 4.3 Baselines
We compare the proposed SDPO with several strong baselines. 1) OpenAI proprietary LLMs. We provide the specific model versions in Appendix D.1. 2) SFT Behavioral Cloning fine-tunes LLMs on expert interaction data, producing a resulting model that serves as the base agent for SDPO and the following baselines. 3) DPO optimizes the agent policy based on data of single turns, specifically targeting the first differing turn in the positive and negative samples used by SDPO. 4) ETO optimizes the agent policy using session-level data. ETO utilizes the same negative sessions as SDPO while sampling five new sessions from scratch to form the data pairs. 5) DMPO leverages the same data as ETO and employs a new loss function to update the policy. 6) Preferred-SFT fine-tunes the base agent on the positive sessions in SDPO.
@@ 4.4 Results
Figure 3: The goal ratings and average words per session for various agents. The word count includes only the utterances of our agents. The square bracket denotes [average turns per session $\times$ average words per turn].
We present the results of SDPO and all the baselines on SOTOPIA in Table 1. As shown, in both goal and relationship dimensions, SDPO significantly outperforms standard DPO, session-level ETO, and DMPO, even surpassing proprietary LLMs like GPT-4o by a large margin, highlighting the effectiveness of segment-level alignment. By analyzing the interaction histories in SOTOPIA, we find that weaker agents often exhibit stubbornness and only express their demands repeatedly. This leads to lower goal and relationship levels, especially in self-chat scenarios. Behavioral cloning using expert data can effectively improve this situation, making the agent more communicative. The reason why Llama-8B+BC’s goal rate drops in its interaction with GPT-4o is that the agent becomes persuadable. We also observe that aligned agents simultaneously improve in both goal and relationship. This indicates that alignment methods indeed enhance the social intelligence of models, rather than achieving goals through behaviors that violate social norms like threatening and deception.
We also repeat the above experiments using Mistral-Instruct-v0.3, with the results presented in Table 2. The detailed experimental setup for Mistral is provided in Appendix D.2. SDPO consistently outperforms all baselines, demonstrating the generalization of our method.
@@ 4.5 Analysis
Variation in Model Output Length We present the output length of various agents during their interactions with GPT-4o in Figure 3. Compared to the BC agent, all alignment methods increase the output length of agents. This phenomenon is commonly observed when DPO is applied to AI chatbots Meng et al. (2024). However, unlike the users’ potential bias toward longer responses, which might be misleading, effective social strategies often require more tokens for communication. Thus, the increase in output length is reasonable. Furthermore, we experiment with terminating the dialogue when the SDPO-tuned agent reaches 10 interaction turns, in order to compare performance under similar token counts. SDPO still outperforms other multi-turn methods, demonstrating that the SDPO-tuned agent utilizes words more efficiently.
Efficiency of Positive Sample Utilization The quality of positive sessions sampled at the session level and segment level is illustrated in Figure 4. When the sampling count is consistent, session-level positives outperform segment-level ones in both goal and relationship. Sampling from scratch provides a larger sampling space compared to using partial interaction history, increasing the likelihood of generating high-quality sessions. However, session-level DPOs, despite using higher-quality data, underperforms SDPO. This indicates that SDPO can more efficiently utilize positive samples due to the finer granularity of segment-level and the theoretically robust loss formulation.
We also analyze the impact of DPO and SDPO on the probability differences between positive and negative samples, as detailed in Appendix E.2.
@@
$[m,n]$ (ETO) 8.19 3.34 7.97 3.01
Table 3: Performance comparison of various segment selection methods. $A_{w}$ denotes the length of the segment selected by GPT-4o from each positive sample, and its distribution is presented in Appendix C.1. $m,n$ represent the sums of all turns in negative and positive samples, respectively, after the erroneous turns.
4.6 Ablation Study
Segment Selection We explore different segment selection methods of SDPO, with the results presented in Table 3. In square brackets, the length of the negative segment is listed first, followed by the positive segment. The segment length refers to the number of turns contained in the segment. For symmetric segment lengths, segments with fixed lengths of 3 and 5 outperform the length of 1 (DPO), demonstrating the efficacy of multi-turn alignment. The segments with length 5 is less effective than that with length 3, indicating that longer segments are not always better. Building on this insight, we leverage GPT-4o to dynamically identify key segments from each positive sample, achieving the best results. For asymmetric segment lengths, model training for segment lengths of [3,1] and [5,3] collapse and can not interact normally. Other asymmetric segments underperform their symmetric counterparts, supporting the theoretical discussions in Section 3.3. Furthermore, we observe that as the degree of asymmetry decreases, the model’s performance improves. This improvement could be attributed to the reduced effect caused by the uneliminated $Z$ on the loss as asymmetry diminishes. This finding helps explain the effectiveness of ETO, which does not impose constraints on the lengths of positive and negative sessions.
@@ 4.6 Ablation Study
Table 4: Performance comparison of models trained on data from different sources. The ‘Source’ column indicates interaction participants: [test agent | interlocutor].
@@ 5 Related Work
Alignment Methods with Refined GranularityRafailov et al. (2023) introduces Direct Preference Optimization (DPO), which utilizes offline data and a streamlined loss function to align LLMs. Various alignment algorithms at refined granularity have been developed based on DPO. Token-level DPO Zeng et al. (2024) integrates forward KL divergence constraints at the token level, enhancing both alignment and diversity. Step-DPO Lai et al. (2024) utilizes individual reasoning steps for preference optimization instead of holistic answer-level evaluation. SePO Yang et al. (2024) presents a token-level reward function estimation method to selectively optimize key tokens in responses. However, in multi-turn interaction scenarios such as social dialogues or web navigation, single-turn alignment is insufficient. To tackle this, ETO and DMPO extend DPO to multi-turn contexts by leveraging session-level data. We take a step further by proposing SDPO, which introduces a dynamic segment-level optimization framework to achieve finer-grained alignment in multi-turn interactions.
@@ 6 Conclusion
In this paper, we introduce Segment-Level Direct Preference Optimization (SDPO) to improve the performance of LLM-based agents in multi-turn social dialogues. Unlike existing multi-turn alignment methods including ETO and DMPO, SDPO focuses on optimizing the agent policy by targeting specific key segments within sessions. Our extensive evaluation on the SOTOPIA benchmark shows that SDPO significantly outperforms existing methods, highlighting the superiority of segment-level alignment. We plan to apply SDPO to other agent tasks to further explore its versatility and validity.
@@ Appendix B Supplementary Theoretical Analysis
B.1 ETO
Song et al. (2024) propose Exploration-Based Trajectory Optimization (ETO), which extends DPO to the session level without rigorous proof. The loss function is as follows:
@@ B.1 ETO
where $h^{w}$, $h^{l}$ represent complete positive and negative interaction histories respectively, $T_{w},T_{l}$ denote the number of rounds in each. When $T_{w}=T_{l}$, the loss function of ETO is equivalent to that of SDPO.
B.2 Discussion on DMPO
Mistake when Applying BT Model After Eq (6), DMPO applies the BT model to obtain the following formula:
$p(\tau^{w}\succ\tau^{l}|h_{0})=\\ \sigma\left(\sum_{t=0}^{T_{w}-1}\gamma^{t}r(h_{t}^{w},y_{t}^{w})-\sum_{t=0}^{T_{l}-1}\gamma^{t}r(h_{t}^{l},y_{t}^{l})\right),$ (10)
where $\tau^{w}$ and $\tau^{l}$ represent the "win" and "lose" samples respectively, $T_{w},T_{l}$ denote the number of rounds in each. A closer examination of Eq (10) reveals that the summation over the $(h,y)$ pairs should exclude $\gamma^{t}$, as it is already incorporated into $d^{\pi}(h,y)$.
In Eq (10), the reward for the entire sequence should be calculated as the summation over all $(h,y)$ pairs. Let’s first discuss why it is valid to sum over time steps $t$. For LLMs, the history $h$ can be viewed as the input context, while $y$ represents the model’s output. In multi-turn interactions, all $(h,y)$ pairs within the sequence are unique. Thus, summing over time steps t is equivalent to summing over $(h,y)$ pairs, making the process more straightforward. However, essentially, each $(h,y)$ pair should be treated equally in the summation, without any inherent concept of time step $t$. Therefore, introducing a discount factor $\gamma^{t}$ is not appropriate.
Limitation of Length Normalization Disregarding the error in Eq. (10) for now, DMPO heuristically introduces regularization for rounds based on Eq. (10) to eliminate $Z$:
$p(\tau^{w}\succ\tau^{l}|h_{0})=\sigma\left(\frac{1-\gamma}{1-\gamma^{T_{w}}}\sum_{t=0}^{T_{w}-1}\gamma^{t}r(h_{t}^{w},y_{t}^{w})\right.\\ \left.-\frac{1-\gamma}{1-\gamma^{T_{l}}}\sum_{t=0}^{T_{l}-1}\gamma^{t}r(h_{t}^{l},y_{t}^{l})\right).$ (11)
However, DMPO did not discuss why length regularization can be applied or the effects brought about by this operation. This transformation lacks rigorous theoretical justification.
DMPO Loss Function Following the DMPO approach, its loss function is as follows:
Figure 6: A detailed overview of three alignment algorithms with varying granularities, illustrated using a SOTOPIA task as an example. Refer to caption represents the agent to be tested. Positive sessions achieve higher scores on the goal and relationship dimensions. For ease of presentation, we simplify the interaction content while preserving the original meaning.
@@ E.1 SOTOPIA Hard Subset
E.2 Necessity of Multi-turn Alignment
After DPO adjusts the first-turn output probabilities for positive and negative segments, will the probabilities of positive segments increase and those of negative segments decrease in subsequent turns? To explore this, we plot the probability differences between positive and negative segments for DPO and SDPO during training, as shown in Figure 9 (only SDPO can be directly compared with DPO; therefore, ETO and DMPO are not mentioned here). The DPO-turn trajectory is nearly parallel to the DPO trajectory, indicating that DPO has almost no influence on the probability differences of subsequent turns. In contrast, the SDPO trajectory rises more steeply. These results demonstrate the necessity of explicitly modifying the probability distribution across turns within the entire segment, providing an explanation for the superiority of multi-turn alignment over DPO.